OpenStudy (anonymous):

2x^3y^4z^3a^2 over 4x^2y^5za Answer choices: A) axz over 2y B) xz^2a over 2y C) ax^2z over 2y D) xy over 2

OpenStudy (kinggeorge):

\[\frac{2x^3y^4z^3a^2}{4x^2y^5za}\]Do you know how to simplify\[\frac{x^3}{x^2}?\]

OpenStudy (anonymous):

no i dont know how to simplify

OpenStudy (anonymous):

@KingGeorge do you know the answer ?

OpenStudy (kinggeorge):

In general, \[\Large \frac{x^a}{x^b}=x^{a-b}\]This means that \[\Large \frac{x^3}{x^2}=x^{3-2}=x^1=x\]Using this information, could you tell me what you think \[\large \frac{y^4}{y^5}\]is?

OpenStudy (anonymous):

wow i have no clue i dont even know how to solve my own problem.